OpenStudy (anonymous):

Can anyone help me with calc 2 material?

OpenStudy (mathteacher1729):

What topics in calc 2?

OpenStudy (anonymous):

I am trying to find a power series representation of\[\arctan(x/\sqrt8)\]

OpenStudy (anonymous):

I would try substitution first: u=x/8^(1/2) then use that derivative of arctan u=1/(1+u^2)=1-u^2+u^4-u^6.... integrate & find constant... helps?

OpenStudy (anonymous):

this may help too: http://www.vias.org/calculus/09_infinite_series_08_04.html

OpenStudy (anonymous):

If i take the derivative of\[f(x)=\arctan(x/\sqrt8)\]I end up with:\[\sqrt8/(8+x^2)\]which can be written\[\sqrt8/8* 1/(1-(-x^2/8))\]this is the power series\[\sqrt8/8*\sum_{n=0}^{\infty}(-1)^n(x^2/8)^n\]if integrate this expression, I should arrive at the equivalent of f(x), since in the first step I took the derivative. Am I correct so far?

OpenStudy (anonymous):

looks good to me! I found this site: http://scidiv.bellevuecollege.edu/dh/ccal/CC10_9.pdf seems usable too

OpenStudy (anonymous):

when I integrate the series I get this:\[\sum_{n=0}^{\infty}(-1)^n((x^2/8)^n(x^2/8))/(n+1)+C\]I am pretty sure this is right, the computer based homework system I am using won't accept this expression though... Any ideas?

OpenStudy (anonymous):

I forgot the factor of sqrt(8)/8 in my last post, still didn't work, any insight would be appreciated

OpenStudy (anonymous):

I 'm looking at something different... see : http://scidiv.bellevuecollege.edu/dh/ccal/CC10_9.pdf p.4 #b) it has arctan( x ) series.. & would just adjust coefficients accordingly, but series should look just like that

OpenStudy (anonymous):

I would assume that the answer should be : \[=\sum_{0}^{\infty}(-1)^{n}* (x/\sqrt{8})^{2n+1}/(2n+1)\] or am I wrong?

OpenStudy (anonymous):

you are right haha i missed that thank you so much!!!

OpenStudy (anonymous):

great! I'm glad we got it.. I was not sure :))